# Organize the Web: Constructing Domains Enhances Pre-Training Data Curation

### v1

@@ Abstract
| Models & Data | huggingface.co/WebOrganizer | |
| Code | github.com/CodeCreator/WebOrganizer |
@@ 1 Introduction
Curating good training data is crucial for enhancing the capabilities of language models. Early pre-training datasets, like the Pile (Gao et al., 2020) or RedPajama (TogetherAI, 2023), were created by curating data from multiple sources—such as Wikipedia, Reddit, or BookCorpus (Zhu et al., 2015)—giving rise to the research problem of how to balance these domains111Throughout the paper, we use the term domain to denote dataset partitions, rather than conventional web domains, which we will refer to as URL domains. (Xie et al., 2023a). However, as the demand for data has grown to trillions of tokens, the majority of data is now obtained from crawling the web, and the importance of curating domains has diminished.
@@ 2 Constructing Domains for Web-Scale Data
State-of-the-art language models rely overwhelmingly on web-crawled training data (Baack, 2024; Raffel et al., 2020; Brown et al., 2020; Rae et al., 2021; Penedo et al., 2023; Soldaini et al., 2024; Penedo et al., 2024; Li et al., 2024)—with open-source research typically resorting to data provided by the CommonCrawl foundation222commoncrawl.org. Unlike the large-scale ImageNet dataset (Deng et al., 2009), which was collected according to an explicit conceptual hierarchy, these web corpora simply contain all web pages adhering to certain filtering rules, resulting in trillions of tokens without an inherent structure. While it is common practice to include selected URL domains, such as Wikipedia or StackOverflow (Touvron et al., 2023; TogetherAI, 2023; Soldaini et al., 2024), these additions are comparatively small and do not demystify the vast amount of data within CommonCrawl.
Our practices of data curation are opaque and uninformed without a firm understanding of how these large-scale corpora are internally composed. In this paper, our approach is to design domain taxonomies to address this short-coming. We first lay out the desirable properties of such domains, then describe our method for creating taxonomies and annotating pre-training datasets, and finally compare our taxonomy-driven domains to a baseline that partitions a corpus via -means clusters.
@@ 2.1 Human-in-the-loop design of domain taxonomies
We design two domain taxonomies for WebOrganizer to capture the topic and format of web pages, respectively. These are meant to capture complementary characteristics: The topic should describe the subject matter of the website content, whereas the format concerns its style, intent and venue.333 This distinction has also been made by van der Wees et al. (2015) in terms of topic and genre.
@@
Our approach of proposing taxonomies in natural language and refining them based on model annotations is flexible and can be used for other purposes—for instance, to annotate a corpus of scientific papers with detailed subject areas, or to taxonomize the data within each of our domains to build a nested hierarchy. Unlike recent techniques for automatically constructing taxonomies with large-language models (Chen et al., 2021; Mishra et al., 2024; Pham et al., 2024), our approach requires human effort, but also benefits from human oversight and domain expertise.
@@ 3 Optimizing Domain Mixtures for Downstream Tasks
Figure 3 visualizes the training distributions predicted by RegMix across the topic and format domains constructed by WebOrganizer. We observe that the two target tasks call for remarkably different data mixtures. The MMLU mixture heavily upsamples Science & Technology, followed by History and Health, and in terms of formats, promotes Academic Writing and Q&A Forums. HellaSwag exhibits smoother mixtures, notably amplifying Home & Hobbies and Fashion & Beauty, and strongly boosting Tutorials. Meanwhile, the mixtures tailored towards the average of both tasks tend to combine the prominent components of each task mixture.
@@ 4.1 Experimental Setting
All our experiments are implemented in the DataComps-LM (DCLM) framework (Li et al., 2024), using the 1b-1x competition pool. We follow best practices and use heuristic filters, followed by deduplication to reduce the 1.6T raw token pool to a base corpus of 200B tokens. From this dataset, we select 29B tokens by sampling according to a domain mixture and train a 1B parameter model. Full details of our experimental setup can be found in Appendix D.
@@ 5 Quality Filters as Implicit Domain Mixers
We adopt only the implicit domain mixtures of quality classifiers for pre-training data curation, replacing the “local” selection within each domain with random sampling. The results of training 1B parameter models are shown in Section 5. Both topic and format domains help to approximate the performance of quality classifiers. Of the two quality classifiers under study, we find FineWeb-Edu to be better approximated by domain effects. In this case, implicit Topic Format mixture recovers its performance gains by 73% on MMLU and 84% on average. However, a substantial gap remains for approximating DCLM-fasttext, suggesting that this classifier relies more on selecting the “right” documents within each domain.
@@
It has become common to claim that datasets which produce better benchmark scores have “higher quality” (Li et al., 2023a; Wettig et al., 2024; Sachdeva et al., 2024; Penedo et al., 2024; Li et al., 2024). In domain mixing, the “quality” of a domain is reflected by how much it should be upsampled, but our findings in Section 3 suggest that MMLU and HellaSwag exhibit very different domain preferences, and optimizing for both tasks requires making trade-offs. This highlights how “data quality” is sensitive to the choice of downstream tasks, and we observe that the notion of “quality” by FineWeb-Edu is particularly biased to specific domains that benefit downstream tasks. However, there are many aspects of web content that are not captured by WebOrganizer, e.g., the prevalence of misspellings or factual errors, and these might be better modeled by scoring individual documents. Such effects may explain why DCLM-fasttext is not well approximated by domain effects, and why both quality filters substantially outperform random sampling when imposing the same Topic Format mixture (Section 5).
@@ 6 Related Work
Many methods have been developed selecting pre-training data for training large language models. It has become common practice to remove noisy web sites using heuristic filtering rules (Raffel et al., 2020; Rae et al., 2021; Penedo et al., 2023), focusing on surface statistics such as mean word length or word repetitions. This is typically followed by deduplication (Lee et al., 2022; Jiang et al., 2023; Abbas et al., 2023; Tirumala et al., 2023; Soldaini et al., 2024). Additional data selection techniques include measuring n-gram similarity to high-quality reference corpora (Brown et al., 2020; Xie et al., 2023b; Li et al., 2024; Brandfonbrener et al., 2024), using perplexity of existing language models (Wenzek et al., 2020; Muennighoff et al., 2023; Marion et al., 2023; Ankner et al., 2025), and prompting large language models to rate documents based on qualities such as factuality or educational value (Gunasekar et al., 2023; Wettig et al., 2024; Sachdeva et al., 2024; Penedo et al., 2024).
@@
Several language models add specially curated domains to their pre-training data (Touvron et al., 2023; Soldaini et al., 2024; OLMo et al., 2024), but CommonCrawl data forms typically the majority of data and has also been shown to outperform domain curation (Penedo et al., 2023; Li et al., 2024). Several works have investigated the specific impact of varying the proportion of code in the pre-training data (Ma et al., 2024; Petty et al., 2024; Aryabumi et al., 2025; Chen et al., 2025). Dubey et al. (2024) briefly mention using knowledge classifiers to downsample “over-represented” data for training Llama-3. Instead of using domains for rebalancing data, Gao et al. (2025) observe performance improvements when conditioning on domain metadata during pre-training.
@@
Many techniques have been developed for tuning domains proportions while training language models. These seek to minimize the validation losses across the domains (Xie et al., 2023a; Albalak et al., 2023; Jiang et al., 2024; Chen et al., 2024), although some methods apply to out-of-domain settings (Chen et al., 2023; Fan et al., 2024). Most methods adjust mixtures dynamically during training, and some make predictions from many static mixtures (Liu et al., 2024; Ye et al., 2024; Kang et al., 2024). Thrush et al. (2025) partition a web corpus into almost 10k domains based on frequent URL domains and rank them based on correlations between domain perplexities and benchmark scores from 90 existing open language models. Hayase et al. (2024) extracts the data mixture of a private pre-training corpus from tokenization rules. Instead of optimizing domain mixtures, researchers have also developed approximations for the impact of training examples on the validation set (Engstrom et al., 2024; Yu et al., 2024; Wang et al., 2024).
@@ 7 Conclusions
Our work advances data curation for language models, and thus carries the broader societal implications associated with improving the capabilities of these models. By taxonomizing web data, we develop a tool that aims to enhance the transparency of pre-training corpora—potentially helping both researchers and the broader public develop a better grasp of the available pre-training data for language models. At the same time, we acknowledge the risks inherent in this process: Reducing the rich diversity of online content to a limited set of discrete domains can obscure important phenomena and may lead to errors, biases, or misrepresentations. We highlight that there are many valid ways to define web taxonomies, and our efforts do not represent a definite “ground truth”. Similarly, the predictions of how to rebalance the domains are sensitive to noise, as they are based on relatively few small model runs. Furthermore, is uncertain how well they transfer across model scales. Despite these challenges, in the absence of other meaningful metadata, we believe that our domain annotations contribute to a more informed understanding of web-scale pre-training data.
@@ Appendix B Training Domain Classifiers
| Topics | Formats | |||
| Avg | Worst | Avg | Worst | |
| Domain classifiers | 93.5 | 87.1 | 91.8 | 80.5 |
| w/o 2-stage training | 91.8 | 84.3 | 90.2 | 74.1 |
| w/o URL features | 92.1 | 86.0 | 88.9 | 80.2 |
@@ Appendix C RegMix Implementation
Appendix D Experimental Details

### v2

@@ Abstract
| Artifacts | hf.co/WebOrganizer | |
| Code | CodeCreator/WebOrganizer |
@@ 1 Introduction
Curating good training data is crucial for enhancing the capabilities of language models. Early pre-training datasets, like the Pile (Gao et al., 2020) or RedPajama (TogetherAI, 2023), were created by curating data from multiple sources—such as Wikipedia, Reddit, or BookCorpus (Zhu et al., 2015)—giving rise to the research problem of how to balance these domains222Throughout the paper, we use the term domain to denote dataset partitions, rather than conventional web domains, which we will refer to as URL domains. (Xie et al., 2023a). However, as the demand for data has grown to trillions of tokens, the majority of data is now obtained from crawling the web, and the importance of curating domains has diminished.
@@ 2 Constructing Domains for Web-Scale Data
State-of-the-art language models rely overwhelmingly on web-crawled training data (Baack, 2024; Raffel et al., 2020; Brown et al., 2020; Rae et al., 2021; Penedo et al., 2023; Soldaini et al., 2024; Penedo et al., 2024; Li et al., 2024)—with open-source research typically resorting to data provided by the CommonCrawl foundation333commoncrawl.org. Unlike the large-scale ImageNet dataset (Deng et al., 2009), which was collected according to an explicit conceptual hierarchy, these web corpora simply contain all web pages adhering to certain filtering rules, resulting in trillions of tokens of text without an inherent structure. While it is common practice to include specially curated domains, such as Wikipedia or StackOverflow (Touvron et al., 2023; TogetherAI, 2023; Soldaini et al., 2024), these additions are comparatively small and do not demystify the vast amount of data within CommonCrawl.
Our practices of data curation are opaque and uninformed without a firm understanding of how these large-scale corpora are internally composed. In this paper, our approach is to design domain taxonomies to address this short-coming. We first lay out the desirable properties of such domains, and then we describe our method for creating taxonomies and annotating pre-training datasets, and finally compare our taxonomy-driven domains to a baseline that partitions a corpus via -means clusters.
@@ 2.1 Human-in-the-loop design of domain taxonomies
We design two domain taxonomies for WebOrganizer to capture the topic and format of web pages, respectively. These are meant to capture complementary characteristics: The topic should describe the subject matter of content, whereas the format concerns its style, intent and venue.444 This distinction has also been made by van der Wees et al. (2015) in terms of topic and genre.
@@
Our approach of proposing taxonomies in natural language and refining them based on model annotations is flexible and can be used for other purposes—for instance, to annotate a corpus of scientific papers with detailed subject areas, or to taxonomize the data within each of our domains to build a nested hierarchy. Unlike recent techniques for automatically constructing taxonomies with large-language models (Chen et al., 2021a; Mishra et al., 2024; Pham et al., 2024), our approach requires human effort, but also benefits from human oversight and domain expertise.
@@ 3 Optimizing Domain Mixtures for Downstream Tasks
Figure 3 visualizes the training distributions predicted by RegMix across the topic and format domains constructed by WebOrganizer. We observe that the two target tasks call for remarkably different data mixtures. The MMLU mixture heavily upsamples Science & Technology, followed by History and Health, and in terms of formats, promotes Academic Writing and Q&A Forums. HellaSwag exhibits smoother mixtures, notably amplifying Home & Hobbies and Fashion & Beauty, and strongly boosting Tutorials. Meanwhile, the mixtures tailored towards the average of both tasks tend to combine the prominent components of each task mixture. We provide predicted mixtures for additional downstream tasks, including HumanEval (Chen et al., 2021b) and Natural Questions (Kwiatkowski et al., 2019) in Appendix D.
@@ 4.1 Experimental Setting
All our experiments are implemented in the DataComps-LM (DCLM) framework (Li et al., 2024), using the 1b-1x competition pool. We follow best practices and use heuristic filters, followed by deduplication to reduce the 1.6T raw token pool to a base corpus of 200B tokens. From this dataset, we select 29B tokens by sampling according to a domain mixture and train a 1B parameter model. Full details of our experimental setup can be found in Appendix E.
@@ 5 Quality Filters as Implicit Domain Mixers
We adopt only the implicit domain mixtures of quality classifiers for pre-training data curation, replacing the “local” selection within each domain with random sampling. The results of training 1B parameter models are shown in Table 2. Both topic and format domains help to approximate the performance of quality classifiers. Of the two quality classifiers under study, we find FineWeb-Edu to be better approximated by domain effects. In this case, implicit Topic Format mixture recovers its performance gains by 73% on MMLU and 84% on average. However, a substantial gap remains for approximating DCLM-fasttext, suggesting that this classifier relies more on selecting the “right” documents within each domain.
@@
It has become common to claim that datasets which produce better benchmark scores have “higher quality” (Li et al., 2023a; Wettig et al., 2024; Sachdeva et al., 2024; Penedo et al., 2024; Li et al., 2024). In domain mixing, the “quality” of a domain is reflected by how much it should be upsampled, but our findings in Section 3 suggest that MMLU and HellaSwag exhibit very different domain preferences, and optimizing for both tasks requires making trade-offs. This highlights how “data quality” is sensitive to the choice of downstream tasks, and we observe that the notion of “quality” by FineWeb-Edu is particularly biased to specific domains that benefit downstream tasks. However, there are many aspects of web content that are not captured by WebOrganizer, e.g., the prevalence of misspellings or factual errors, and these might be better modeled by scoring individual documents. Such effects may explain why DCLM-fasttext is not well approximated by domain effects, and why both quality filters substantially outperform random sampling when imposing the same Topic Format mixture (Table 2).
@@ 6 Related Work
Many methods have been developed selecting pre-training data for training large language models. It has become common practice to remove noisy web sites using heuristic filtering rules (Raffel et al., 2020; Rae et al., 2021; Penedo et al., 2023), focusing on surface statistics such as mean word length or word repetitions. This is typically followed by deduplication (Lee et al., 2022; Jiang et al., 2023; Abbas et al., 2023; Tirumala et al., 2023; Soldaini et al., 2024). Additional data selection techniques include measuring n-gram similarity to high-quality reference corpora (Brown et al., 2020; Xie et al., 2023b; Li et al., 2024; Brandfonbrener et al., 2024), using perplexity of existing language models (Wenzek et al., 2020; Muennighoff et al., 2023; Marion et al., 2023; Ankner et al., 2025), and prompting large language models to rate documents based on qualities such as factuality or educational value (Gunasekar et al., 2023; Wettig et al., 2024; Sachdeva et al., 2024; Penedo et al., 2024). An alternative approach to data curation focuses on generating synthetic training data from a large language models (Gunasekar et al., 2023; Li et al., 2023a), which may involve designing an explicit taxonomy of skills and concepts (Ding et al., 2023; Ben Allal et al., 2024) or re-writing existing data (Maini et al., 2024)—the latter would reflect the original domain proportions and our approach could be used to improve the data mixture.
@@
Several language models add specially curated domains to their pre-training data (Touvron et al., 2023; Soldaini et al., 2024; OLMo et al., 2024), but CommonCrawl data forms typically the majority of data and has also been shown to outperform domain curation (Penedo et al., 2023; Li et al., 2024). Several works have investigated the specific impact of varying the proportion of code in the pre-training data (Ma et al., 2024; Petty et al., 2024; Aryabumi et al., 2025; Chen et al., 2025). Dubey et al. (2024) briefly mention using knowledge classifiers to downsample “over-represented” data for training Llama-3. Instead of using domains for rebalancing data, Gao et al. (2025) observe performance improvements when conditioning on domain metadata during pre-training. Similar to our work, (Bai et al., 2024) propose to classify data into 13 topics, and combine this with FineWeb-Edu quality buckets and SlimPajama sources, resulting in a large set of composite domains. Rather than learning the relationship between these domain weights and downstream tasks with RegMix, they propose a selection strategy that samples from the domains based on gradient influence scores. Similarly, (Zhang et al., 2025) define fine-grained -means clusters (=10,000) with the motivation to increase diversity during influence-based data selection. SemDeDup (Abbas et al., 2023) also utilizes a large set of -means clusters to define prototypicality scores and diversify the pre-training distribution. While all these works have the shared goal of improving data quality, our work also contributes a two-dimensional structuree for organizing web data and provides a comparative study of fine-grained quality selection and coarse-grained domain mixing.
@@
Many techniques have been developed for tuning domains proportions while training language models. These seek to minimize the validation losses across the domains (Xie et al., 2023a; Albalak et al., 2023; Jiang et al., 2024; Chen et al., 2024), although some methods apply to out-of-domain settings (Chen et al., 2023; Fan et al., 2024). Most methods adjust mixtures dynamically during training, and some make predictions from many static mixtures (Liu et al., 2024; Ye et al., 2024; Kang et al., 2024). In concurrent work, Held et al. (2025) use large language model to predict the utility of subsets to downstream tasks. Due to the lack of meaningful domains, CommonCrawl is partitioned into “head” and “tail” domains based on perplexity scores. Thrush et al. (2025) partition a web corpus into almost 10k domains based on frequent URL domains and rank them based on correlations between domain perplexities and benchmark scores from 90 existing open language models. Hayase et al. (2024) extracts the data mixture of a private pre-training corpus from tokenization rules. Instead of optimizing domain mixtures, researchers have also developed approximations for the impact of individual training examples on loss of a validation set (Engstrom et al., 2024; Yu et al., 2024; Wang et al., 2024). However, (Zhang et al., 2025) demonstrate that partitioning the dataset into domains (in their work, -means clusters) is beneficial for increasing data diversity when selecting data with gradient-based influence approximations.
@@ 7 Conclusions
Our work advances data curation for language models, and thus carries the broader societal implications associated with improving the capabilities of these models. By taxonomizing web data, we develop a tool that aims to enhance the transparency of pre-training corpora—potentially helping both researchers and the broader public develop a better grasp of the available pre-training data for language models. At the same time, we acknowledge the risks inherent in this process: Reducing the rich diversity of online content to a limited set of discrete domains can obscure important phenomena and may lead to errors, biases, or misrepresentations. We highlight that there are many valid ways to define web taxonomies, and our efforts do not represent a definite “ground truth”. Similarly, the predictions of how to rebalance the domains are sensitive to noise, as they are based on relatively few small model runs. Furthermore, is uncertain how well they transfer across model scales. Despite these challenges, in the absence of other meaningful metadata, we believe that our domain annotations contribute to a more informed understanding of web-scale training data.
@@ Appendix C RegMix Implementation
| Topics | Formats | |||
| Avg | Worst | Avg | Worst | |
| Domain classifiers | 93.5 | 87.1 | 91.8 | 80.5 |
| w/o 2-stage training | 91.8 | 84.3 | 90.2 | 74.1 |
| w/o URL features | 92.1 | 86.0 | 88.9 | 80.2 |
@@
Appendix D Predicted Mixtures
Table 9 reports the numerical results of the mixtures visualized in Figure 3. In addition to our main two tasks of focus, we also include predicted data mixtures for a wider range of downstream tasks in Figure 8. Note that we use bits-per-bytes of the correct solution across all tasks. We use 5 in-context examples for MMLU, HellaSwag, Natural Questions (NQ), 3 examples for HumanEval and MBPP, 0-shot for MATH. We note that both coding tasks upsample documents from the Software Engineering topic and Documentation format. Natural Questions also exhibits distinct patterns as being the only task to upsample the Entertainment topic category heavily.
@@ Appendix D Predicted Mixtures
Appendix E Experimental Details